cpl106:

A fair, unbiased coin was flipped 10 times, giving the results shown in the table, where T = tails and H = heads. Flip 1 2 3 4 5 6 7 8 9 10 Result T T T H T T T H T T What is the difference between the theoretical and experimental probabilities of getting heads? A. 0.5 B. 0.0 C. 0.1 D. 0.3

cpl106:

jimthompson5910:

How many times does heads show up in the table?

cpl106:

8

jimthompson5910:

check again

jimthompson5910:

count how many times "H" shows up

cpl106:

2

jimthompson5910:

yes

cpl106:

i read it wrong mu bad

jimthompson5910:

so we divide that over the total number of flips (10) that will give us the experimental probability (aka empirical probability)

jimthompson5910:

what do you get for the experimental probability?

cpl106:

0.5

jimthompson5910:

A = number of heads in the table = 2 B = number of flips = 10 A/B = ?

cpl106:

5?

jimthompson5910:

2/10 = ?

jimthompson5910:

you might be doing 10/2 which is the wrong order

cpl106:

0.2

jimthompson5910:

yes

jimthompson5910:

Now for the theoretical probability 1 side is heads and there are 2 sides 1/2 = 0.5 is the theoretical probability of getting heads

jimthompson5910:

the last thing you do is subtract the theoretical and experimental probabilities

cpl106:

0.3?

jimthompson5910:

yes

cpl106:

ty

jimthompson5910:

no problem